Circular No. 6977

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998de IN NGC 252 M. Modjaz, E. Halderson, T. Shefler, J. Y. King, W. D. Li, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998de was found and confirmed in unfiltered CCD images obtained on July 23.5 (mag ∼ 18.4) and 24.5 UT (mag ∼ 17.8). The object is located at α = 0^{h}48^{m}06^{s}.88, δ = +27^{o}37′28′′.5 (equinox 2000.0), which is ∼ 71′′.9 east and 3′′.4 north of the nucleus of NGC 252. A KAIT image on July 19.5 already showed a hint of the supernova, and a KAIT image on July 14.5 did not show any star at the supernova’s position (limiting mag ∼ 19.0).

COMET C/1998 H1 (STONEHOUSE)

Ephemeris extension to IAUC 6930 (elements MPC 31893):

1998 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 16 12^{h}16.^{m}96 +50^{o}45.^{′}9 2.247 1.956 60.^{o}5 26.^{o}9 14.4
26 12 23.08 +50 43.7 2.419 2.044 56.5 24.5 14.7
Aug. 5 12 31.35 +50 45.2 2.571 2.134 53.7 22.5 15.0
15 12 41.39 +50 52.5 2.702 2.227 52.0 21.0 15.3
25 12 53.02 +51 07.2 2.811 2.321 51.5 19.9 15.5
Sept. 4 13 06.11 +51 31.0 2.901 2.418 52.2 19.2 15.8
14 13 20.61 +52 05.2 2.972 2.515 54.0 18.9 16.0
24 13 36.56 +52 51.0 3.028 2.614 56.6 18.7 16.2

Total visual magnitude estimates: July 18.88 UT, 14.6 (M. Lehký, Hradec Králové, Czech Republic, 0.42-m reflector; 20.88, 14.5 (Lehký); 20.90, 13.9 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector; 21.89, 14.0 (Hornoch).

COMET C/1998 K5 (LINEAR) Visual m_{1} estimates (L = reflector): June 18.27 UT, 13.2 (A. Hale,

Cloudcroft, NM, 0.41-m L); 25.97, 12.9 (K. Sárneczky, Ágasvár, Hungary, 0.44-m L); July 2.89, 12.1 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m L); 19.98, 12.5 (M. Meyer, Frauenstein, Germany, 0.25-m L); 23.01, 12.5 (R. J. Bouma, Lattrop, The Netherlands, 0.40-m L).

1998 July 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green